# Anisotropy of quadratic forms over global fields of characteristic $\neq$ 2 is Diophantine

### v2

@@ Abstract.
We prove that the set of anisotropic quadratic forms over global fields of characteristic different from 2 is a diophantine set. Our proof builds upon and extends the method of Koenigsmann, using tools from class field theory, local-global principles, and recent advances on the diophantine definability of non-norm sets over global fields.
@@ 1. Introduction
Let $K$ be a global field. Roughly speaking, a diophantine set is a set that can be described by the solvability of a family of polynomial equations. Equivalently, we may give the following definition.
@@
This result has been generalized to various situations in recent years. Colliot-Thélène and Van Geel [CTG15] extended Poonen’s result to the complements of $n$-th powers over global fields. Koenigsmann [Koe16] introduced powerful new methods for constructing diophantine sets over $\mathbb{Q}$, notably the set of non-norms for quadratic extensions.
@@
In this paper, we turn to diophantine sets defined by quadratic forms. Our goal is to show that the set of coefficients of anisotropic quadratic forms over a global field of characteristic not equal to $2$ is a diophantine set. In any case, the set of coefficients of isotropic quadratic forms is obviously diophantine. So, the set we deal with here is the complement of this obvious diophantine set, just like the works we mentioned above. Our main theorem is the following.
@@
In the notation $(K^{\times})^{m}$ we mean the set of $m$-tuples of elements of $K^{\times}$, not the set of $m$th powers in $K^{\times}$, which we denote by ${K^{\times}}^{m}$.
We say that a quadratic form $f$ represents 0 over $K$ if there exists a non-trivial tuple $(x_{1},\dots,x_{m})\in K^{m}\setminus\{0\}$ such that $f(x_{1},\dots,x_{m})=0$.
As a corollary, we obtain the following:
@@
$\bigl\{(a_{ij})\in(K^{\times})^{m^{2}}:\sum_{1\leq i,j\leq m}a_{ij}x_{i}x_{j}\text{ is anisotropic over }K\bigr\}$
forms a diophantine set over $K$.
To prove Theorem 1.2, we consider different values of $m$. For $m\geq 5$, the result follows from local-global principles. For $m=1,2,$ or $3$, the result is already known in the context of “non-squares” and “non-norms”. The only genuinely new case is $m=4$, for which we employ generalizations of Koenigsmann’s method as developed in [Dit18, Par13, EM18].
@@ 2. Preliminaries
From now on, we use $S_{K}$ to denote the set of primes of $K$, including both finite and infinite primes. By a prime of $K$ we mean an equivalence class of nontrivial absolute values of $K$.
For a number field $K$, we write $S_{K}^{\infty}$ for the set of infinite primes. For $\mathfrak{p}\in S_{K}^{\infty}$, if the corresponding completion $K_{\mathfrak{p}}$ is isomorphic to the real field $\mathbb{R}$, we call $\mathfrak{p}$ a real infinite prime; if $K_{\mathfrak{p}}\simeq\mathbb{C}$, then $\mathfrak{p}$ is called a complex infinite prime.
[…]
For $a,b\in K_{\mathfrak{p}}^{\times}$, the notation $(a,b)_{\mathfrak{p}}$ denotes the Hilbert symbol at $\mathfrak{p}$. Note that the Hilbert symbol $(a,b)_{\mathfrak{p}}$ is a nondegenerate bilinear pairing on $K_{\mathfrak{p}}^{\times}/{K_{\mathfrak{p}}^{\times}}^{2}$. For $a\in K_{\mathfrak{p}}^{\times}/{K_{\mathfrak{p}}^{\times}}^{2}$, we define the homomorphism of groups
$T_{a}:K_{\mathfrak{p}}^{\times}/{K_{\mathfrak{p}}^{\times}}^{2}\rightarrow\{\pm 1\}$
@@
$x\mapsto(x,a)_{\mathfrak{p}},$
where we regard $\{\pm 1\}$ as a group by multiplication.
@@
With the notations above, for $\epsilon,\epsilon^{\prime}\in\{\pm 1\}$ and nontrivial $a,a^{\prime}\in K_{\mathfrak{p}}^{\times}/{K_{\mathfrak{p}}^{\times}}^{2}$, we have $T_{a}^{-1}(\epsilon)\cap T^{-1}_{a^{\prime}}(\epsilon^{\prime})=\emptyset\Leftrightarrow a=a^{\prime}\wedge\epsilon=-\epsilon^{\prime}$.
@@
$\Rightarrow:$ By 63:9 of [O’M00] we know that the index $K_{\mathfrak{p}}^{\times}:{K_{\mathfrak{p}}^{\times}}^{2}]$ is finite. $T_{a}$ is nontrivial since $a\neq 1$ and the Hilbert symbol is nondegenerate, hence the cardinality of $T_{a}^{-1}(1)$ and $T_{a}^{-1}(-1)$ are both half of $K_{\mathfrak{p}}^{\times}:{K_{\mathfrak{p}}^{\times}}^{2}]$.
For $\epsilon,\epsilon^{\prime}\in\{\pm 1\}$ and nontrivial $a,a^{\prime}\in K_{\mathfrak{p}}^{\times}/{K_{\mathfrak{p}}^{\times}}^{2}$, we have
$T_{a}^{-1}(\epsilon)\cap T^{-1}_{a^{\prime}}(\epsilon^{\prime})=\emptyset\Leftrightarrow T_{a}^{-1}(\epsilon)\sqcup T^{-1}_{a^{\prime}}(\epsilon^{\prime})=K_{\mathfrak{p}}^{\times}/{K_{\mathfrak{p}}^{\times}}^{2}$
by the reason of cardinality.
The complement of $T_{a}^{-1}(1)$ in $K_{\mathfrak{p}}^{\times}/{K_{\mathfrak{p}}^{\times}}^{2}$ is $T_{a}^{-1}(-1)$, so
$T_{a}^{-1}(\epsilon)\sqcup T^{-1}_{a^{\prime}}(\epsilon^{\prime})=K_{\mathfrak{p}}^{\times}/{K_{\mathfrak{p}}^{\times}}^{2}$
is equivalent to
[…]
However, $T_{a}^{-1}(1)=T_{a^{\prime}}^{-1}(-1)$ is impossible, since this will imply
$(x,aa^{\prime})_{\mathfrak{p}}=(x,a)_{\mathfrak{p}}(x,a^{\prime})_{\mathfrak{p}}=-1,\forall x\in K_{\mathfrak{p}}^{\times}/{K_{\mathfrak{p}}^{\times}}^{2},$
@@
$T_{a}^{-1}(1)=T_{a^{\prime}}^{-1}(1)$ $\Leftrightarrow(x,a)_{\mathfrak{p}}=(x,a^{\prime})_{\mathfrak{p}},\forall x\in K_{\mathfrak{p}}^{\times}/{K_{\mathfrak{p}}^{\times}}^{2}$
$\Leftrightarrow(x,aa^{\prime})_{\mathfrak{p}}=1,\forall x\in K_{\mathfrak{p}}^{\times}/{K_{\mathfrak{p}}^{\times}}^{2}$
$\Leftrightarrow T_{aa^{\prime}}\textup{ is trivial}$
$\Leftrightarrow aa^{\prime}=1$
$\Leftrightarrow a=a^{\prime}.$
In conclusion, we have $T_{a}^{-1}(\epsilon)\cap T^{-1}_{a^{\prime}}(\epsilon^{\prime})=\emptyset\Leftrightarrow a=a^{\prime}\wedge\epsilon=-\epsilon^{\prime}$. ∎
@@ 2.1. Quadratic forms
The arithmetic of quadratic forms over a global field is well understood. In the case of fields with characteristic not equal to $2$, the situation is even simpler, as the local-global principle applies.
For proving the Theorem 2.3 below, we need the following lemma:
@@
[…]
For (1) $\Rightarrow$ (2), suppose that there exists
[…]
@@
If $a_{1}x_{1}^{2}+a_{2}x_{2}^{2}\neq 0$, take $x^{\prime}=a_{1}x_{1}^{2}+a_{2}x_{2}^{2}$ then we are done. If $a_{1}x_{1}^{2}+a_{2}x_{2}^{2}=0$, then the quadratic form $a_{1}x_{1}^{2}+a_{2}x_{2}^{2}$ is isotropic, hence universal by 42:10 of [O’M00]. Then we can take $x^{\prime}=-a_{3}x_{3}^{2}-a_{4}x_{4}^{2}$ for arbitrary $(x_{3},x_{4})\in(K_{\mathfrak{p}})^{2}$ such that $x^{\prime}\neq 0$. ∎
@@
    then $f$ is anisotropic over $K_{\mathfrak{p}}$ if and only if $a_{1}a_{2}a_{3}a_{4}\in{K_{\mathfrak{p}}}^{\times 2}$ and $(a_{1},a_{2})_{\mathfrak{p}}=-(-a_{3},-a_{4})_{\mathfrak{p}}$.
@@
The first item follows from Theorem 66:1 in [O’M00], and the second item from Theorem 63:19 in [O’M00]. The third item is stated in Theorem 6 of Chapter IV in [Ser73] for $K=\mathbb{Q}$. The same proof can be naturally generalized to global fields. We provide a detailed proof of this generalization below, following the method in [Ser73].
By Lemma 2.2, $f=a_{1}x_{1}^{2}+a_{2}x_{2}^{2}+a_{3}x_{3}^{2}+a_{4}x_{4}^{2}$ is isotropic over $K_{\mathfrak{p}}$ if and only if there exists a nonerzo $x^{\prime}\in K_{\mathfrak{p}}$ such that the two forms $a_{1}x_{1}^{2}+a_{2}x_{2}^{2}$ and $a_{3}x_{3}^{2}-a_{4}x_{4}^{2}$ represents $x^{\prime}$. Let $x$ denote the coset of $x^{\prime}$ in $K_{\mathfrak{p}}^{\times}/{K_{\mathfrak{p}}^{\times}}^{2}$. The form $a_{1}x_{1}^{2}+a_{2}x_{2}^{2}$ represents $x^{\prime}$ is equivalent to $(xa_{1},xa_{2})_{\mathfrak{p}}=1$. After a calculation of Hilbert symbol, we get:
$(xa_{1},xa_{2})_{\mathfrak{p}}$ $=(xa_{1},x)_{\mathfrak{p}}(xa_{1},a_{2})_{\mathfrak{p}}$
$=(x,x)_{\mathfrak{p}}(a_{1},x)_{\mathfrak{p}}(x,a_{2})_{\mathfrak{p}}(a_{1},a_{2})_{\mathfrak{p}}$
$=(x,-1)_{\mathfrak{p}}(x,a_{1})_{\mathfrak{p}}(x,a_{2})_{\mathfrak{p}}(a_{1},a_{2})_{\mathfrak{p}}$
$=(x,-a_{1}a_{2})_{\mathfrak{p}}(a_{1},a_{2})_{\mathfrak{p}}.$
So $a_{1}x_{1}^{2}+a_{2}x_{2}^{2}$ represents $x^{\prime}$ is equivalent to
@@
which means exactly $x\in T_{-a_{1}a_{2}}^{-1}((a_{1},a_{2})_{\mathfrak{p}})$.
For the same reason, $a_{3}x_{1}^{2}-a_{4}x_{2}^{2}$ represents $x^{\prime}$ is equivalent to
@@
which means exactly $x\in T_{-a_{3}a_{4}}^{-1}((-a_{3},-a_{4})_{\mathfrak{p}})$.
Hence the absence of such an $x$ is equivalent to $T_{-a_{1}a_{2}}^{-1}((a_{1},a_{2})_{\mathfrak{p}})\cap T_{-a_{3}a_{4}}^{-1}((-a_{3},-a_{4})_{\mathfrak{p}})=\emptyset$. By Lemma 2.1, this is equivalent to
[…]

### v3

# arXiv:2401.00537v3[math.NT] 28 Feb 2026

@@ Abstract.
We prove that the set of anisotropic quadratic forms over global fields of characteristic different from 2 is a diophantine set. Our proof builds upon and extends the method of Koenigsmann, using tools from class field theory, the local-global principle, and advances on the diophantine definability of non-norm sets over global fields.
@@ 1. Introduction
Let $K$ be a global field. Roughly speaking, a diophantine set is a set that can be described by the solvability of a family of polynomial equations. Equivalently, we give the following definition.
@@
This result has been generalized to various situations in recent years. Colliot-Thélène and Jan Van Geel [CG15] extended Poonen’s result to the complements of $n$-th powers over global fields. Koenigsmann [KOE16] introduced powerful new methods for constructing diophantine sets over $\mathbb{Q}$, notably the set of non-norms for quadratic extensions.
@@
In this paper, we study diophantine sets defined by quadratic forms. We begin by recalling the notions of isotropy and anisotropy of quadratic forms.
Definition 1.2.
Let $F$ be a field and let $f$ be a quadratic form in $m$ variables over $F$. We say that $f$ is isotropic over $F$ if there exists a tuple $(x_{1},\dots,x_{m})\in F^{m}$, not all zero, such that
$f(x_{1},\dots,x_{m})=0.$
Otherwise, we say that $f$ is anisotropic over $F$.
Our goal is to show that the set of coefficients of anisotropic quadratic forms over a global field of characteristic not equal to $2$ is a diophantine set. In any case, the set of coefficients of isotropic quadratic forms is obviously diophantine. So, the set we deal with here is the complement of this obvious diophantine set, just like the works we mentioned above. Our main theorem is the following.
@@
In the notation $(K^{\times})^{m}$ we mean the Cartesian product of $m$ copies of $K^{\times}$, i.e. the set of $m$-tuples of elements of $K^{\times}$, and not the set of $m$th powers in $K^{\times}$, which we denote by ${K^{\times}}^{m}$.
As a corollary of Theorem 1.3, we obtain the following.
@@
Let $K$ be a global field with $\mathrm{char}(K)\neq 2$, and let $m\in\mathbb{N}$. Then the set
$\bigl\{(a_{ij})\in(K)^{m^{2}}\;\big|\;\sum_{1\leq i,j\leq m}a_{ij}x_{i}x_{j}\text{ is anisotropic over }K\bigr\}$
is a diophantine set over $K$.
To prove Theorem 1.3, we distinguish cases according to the value of $m$. For $m\geq 5$, the result follows from local–global principles. For $m=1,2,$ or $3$, the result is already known, in the context of “non-squares” and “non-norms”. The only genuinely new case is $m=4$, for which we employ generalizations of Koenigsmann’s method developed in [DIT18, PAR13, EM18].
@@ 2. Preliminaries
From now on, we denote by $S_{K}$ the set of primes of $K$, including both finite and infinite primes. By a prime of $K$ we mean an equivalence class of nontrivial absolute values on $K$.
If $K$ is a number field, we write $S_{K}^{\infty}$ for the set of infinite primes of $K$. For $\mathfrak{p}\in S_{K}^{\infty}$, if the corresponding completion $K_{\mathfrak{p}}$ is isomorphic to the real field $\mathbb{R}$, we call $\mathfrak{p}$ a real infinite prime; if $K_{\mathfrak{p}}\simeq\mathbb{C}$, then $\mathfrak{p}$ is called a complex infinite prime. If $\mathfrak{p}$ is finite, we also use $\mathfrak{p}$ to denote the corresponding prime ideal, and we write $v_{\mathfrak{p}}$ for the associated normalized discrete valuation on $K$. Let
$S_{K}^{f}\coloneqq S_{K}\setminus S_{K}^{\infty}$
denote the set of finite primes of $K$.
For $a,b\in K_{\mathfrak{p}}^{\times}$, the notation $(a,b)_{\mathfrak{p}}$ denotes the Hilbert symbol at $\mathfrak{p}$. Note that the Hilbert symbol $(a,b)_{\mathfrak{p}}$ defines a nondegenerate bilinear pairing on $K_{\mathfrak{p}}^{\times}/K_{\mathfrak{p}}^{\times 2}$.
For $a\in K_{\mathfrak{p}}^{\times}/K_{\mathfrak{p}}^{\times 2}$, we define the group homomorphism
$T_{a}\colon K_{\mathfrak{p}}^{\times}/K_{\mathfrak{p}}^{\times 2}\longrightarrow\{\pm 1\}$
@@
$x\longmapsto(x,a)_{\mathfrak{p}},$
where we regard $\{\pm 1\}$ as a multiplicative group. Let $f$ be a quadratic form in $m$ variables over $K$, and let $a\in K^{\times}$. We say that $f$ represents $a$ over $K$ if there exists a tuple $(x_{1},\dots,x_{m})\in K^{m}$ such that
$f(x_{1},\dots,x_{m})=a.$
We say that $f$ is isotropic (or anisotropic) over $\mathfrak{p}$ if it is isotropic (or anisotropic) over the completion $K_{\mathfrak{p}}$.
@@
With the notations above, let $\epsilon,\epsilon^{\prime}\in\{\pm 1\}$ and let $a,a^{\prime}\in K_{\mathfrak{p}}^{\times}/K_{\mathfrak{p}}^{\times 2}$ be nontrivial. Then
$T_{a}^{-1}(\epsilon)\cap T_{a^{\prime}}^{-1}(\epsilon^{\prime})=\emptyset\quad\Longleftrightarrow\quad a=a^{\prime}\text{ and }\epsilon=-\epsilon^{\prime}.$
@@
$\Rightarrow:$ By 63:9 of [O’M00], the index $K_{\mathfrak{p}}^{\times}:K_{\mathfrak{p}}^{\times 2}]$ is finite. Since $a\neq 1$ and the Hilbert symbol is nondegenerate, $T_{a}$ is nontrivial. Hence, the cardinalities of $T_{a}^{-1}(1)$ and $T_{a}^{-1}(-1)$ are both equal to half of $K_{\mathfrak{p}}^{\times}:K_{\mathfrak{p}}^{\times 2}]$.
For $\epsilon,\epsilon^{\prime}\in\{\pm 1\}$ and nontrivial $a,a^{\prime}\in K_{\mathfrak{p}}^{\times}/K_{\mathfrak{p}}^{\times 2}$, we have
$T_{a}^{-1}(\epsilon)\cap T_{a^{\prime}}^{-1}(\epsilon^{\prime})=\emptyset\quad\Longleftrightarrow\quad T_{a}^{-1}(\epsilon)\sqcup T_{a^{\prime}}^{-1}(\epsilon^{\prime})=K_{\mathfrak{p}}^{\times}/K_{\mathfrak{p}}^{\times 2},$
by a cardinality argument.
The complement of $T_{a}^{-1}(1)$ in $K_{\mathfrak{p}}^{\times}/K_{\mathfrak{p}}^{\times 2}$ is $T_{a}^{-1}(-1)$. Hence,
$T_{a}^{-1}(\epsilon)\sqcup T_{a^{\prime}}^{-1}(\epsilon^{\prime})=K_{\mathfrak{p}}^{\times}/K_{\mathfrak{p}}^{\times 2}\quad\text{is equivalent to}\quad T_{a}^{-1}(1)=T_{a^{\prime}}^{-1}(1)\text{ or }T_{a}^{-1}(1)=T_{a^{\prime}}^{-1}(-1).$
However, $T_{a}^{-1}(1)=T_{a^{\prime}}^{-1}(-1)$ is impossible, since this would imply
$(x,aa^{\prime})_{\mathfrak{p}}=(x,a)_{\mathfrak{p}}(x,a^{\prime})_{\mathfrak{p}}=-1\quad\forall x\in K_{\mathfrak{p}}^{\times}/K_{\mathfrak{p}}^{\times 2},$
@@
$T_{a}^{-1}(1)=T_{a^{\prime}}^{-1}(1)$ $\Longleftrightarrow(x,a)_{\mathfrak{p}}=(x,a^{\prime})_{\mathfrak{p}}\quad\forall x\in K_{\mathfrak{p}}^{\times}/K_{\mathfrak{p}}^{\times 2}$
$\Longleftrightarrow(x,aa^{\prime})_{\mathfrak{p}}=1\quad\forall x\in K_{\mathfrak{p}}^{\times}/K_{\mathfrak{p}}^{\times 2}$
$\Longleftrightarrow T_{aa^{\prime}}\text{ is trivial}$
$\Longleftrightarrow aa^{\prime}=1$
$\Longleftrightarrow a=a^{\prime}.$
In conclusion, we have
$T_{a}^{-1}(\epsilon)\cap T_{a^{\prime}}^{-1}(\epsilon^{\prime})=\emptyset\quad\Longleftrightarrow\quad a=a^{\prime}\text{ and }\epsilon=-\epsilon^{\prime}.$
@@ 2.1. Quadratic forms
The arithmetic of quadratic forms over a global field is well understood. When the characteristic of the field is not $2$, the situation simplifies further, as the local-global principle applies.
To prove Theorem 2.3 below, we will need the following lemma:
@@
(2) $\Rightarrow$ (1) is obvious.
For (1) $\Rightarrow$ (2), suppose there exists
$(x_{1},x_{2},x_{3},x_{4})\in(K_{\mathfrak{p}})^{4}\setminus\{(0,0,0,0)\}$
@@
If $a_{1}x_{1}^{2}+a_{2}x_{2}^{2}\neq 0$, take $x^{\prime}=a_{1}x_{1}^{2}+a_{2}x_{2}^{2}$, and we are done. If $a_{1}x_{1}^{2}+a_{2}x_{2}^{2}=0$, then the quadratic form $a_{1}x_{1}^{2}+a_{2}x_{2}^{2}$ is isotropic, and hence universal by 42:10 of [O’M00]. In this case, we can take $x^{\prime}=-a_{3}x_{3}^{2}-a_{4}x_{4}^{2}$ for any $(x_{3},x_{4})\in(K_{\mathfrak{p}})^{2}$ such that $x^{\prime}\neq 0$. ∎
@@
    then $f$ is anisotropic over $K_{\mathfrak{p}}$ if and only if
    $a_{1}a_{2}a_{3}a_{4}\in K_{\mathfrak{p}}^{\times 2}\quad\text{and}\quad(a_{1},a_{2})_{\mathfrak{p}}=-(-a_{3},-a_{4})_{\mathfrak{p}}.$
@@
The first item follows from Theorem 66:1 in [O’M00], and the second item from Theorem 63:19 in [O’M00]. The third item is stated in Theorem 6 of Chapter IV in [SER73] for $K=\mathbb{Q}$, and the same proof naturally generalizes to arbitrary global fields. We provide a detailed proof of this generalization below, following the method in [SER73].
By Lemma 2.2, the form $f=a_{1}x_{1}^{2}+a_{2}x_{2}^{2}+a_{3}x_{3}^{2}+a_{4}x_{4}^{2}$ is isotropic over $K_{\mathfrak{p}}$ if and only if there exists a nonzero $x^{\prime}\in K_{\mathfrak{p}}$ such that the two forms $a_{1}x_{1}^{2}+a_{2}x_{2}^{2}$ and $a_{3}x_{3}^{2}-a_{4}x_{4}^{2}$ both represent $x^{\prime}$. Let $x$ denote the coset of $x^{\prime}$ in $K_{\mathfrak{p}}^{\times}/K_{\mathfrak{p}}^{\times 2}$.
The statement that $a_{1}x_{1}^{2}+a_{2}x_{2}^{2}$ represents $x^{\prime}$ is equivalent to
$(xa_{1},xa_{2})_{\mathfrak{p}}=1.$
After a calculation using the properties of the Hilbert symbol, we get
$(xa_{1},xa_{2})_{\mathfrak{p}}$ $=(xa_{1},x)_{\mathfrak{p}},(xa_{1},a_{2})_{\mathfrak{p}}$
$=(x,x)_{\mathfrak{p}},(a_{1},x)_{\mathfrak{p}},(x,a_{2})_{\mathfrak{p}},(a_{1},a_{2})_{\mathfrak{p}}$
$=(x,-1)_{\mathfrak{p}},(x,a_{1})_{\mathfrak{p}},(x,a_{2})_{\mathfrak{p}},(a_{1},a_{2})_{\mathfrak{p}}$
$=(x,-a_{1}a_{2})_{\mathfrak{p}},(a_{1},a_{2})_{\mathfrak{p}}.$
Hence, $a_{1}x_{1}^{2}+a_{2}x_{2}^{2}$ represents $x^{\prime}$ if and only if
@@
which is equivalent to
$x\in T_{-a_{1}a_{2}}^{-1}((a_{1},a_{2})_{\mathfrak{p}}).$
Similarly, $a_{3}x_{3}^{2}-a_{4}x_{4}^{2}$ represents $x^{\prime}$ if and only if
@@
which is equivalent to
$x\in T_{-a_{3}a_{4}}^{-1}((-a_{3},-a_{4})_{\mathfrak{p}}).$
Therefore, the absence of such an $x$ is equivalent to
$T_{-a_{1}a_{2}}^{-1}((a_{1},a_{2})_{\mathfrak{p}})\cap T_{-a_{3}a_{4}}^{-1}((-a_{3},-a_{4})_{\mathfrak{p}})=\emptyset.$
By Lemma 2.1, this is in turn equivalent to
$a_{1}a_{2}=a_{3}a_{4}\text{ up to a square in }K_{\mathfrak{p}}^{\times}\quad\text{and}\quad(a_{1},a_{2})_{\mathfrak{p}}=-(-a_{3},-a_{4})_{\mathfrak{p}},$
as claimed. ∎